### v1

# Safe-Control: A Safety Patch for Mitigating Unsafe Content in Text-to-Image Generation Models

Abstract
Despite the advancements in Text-to-Image (T2I) generation models, their potential for misuse or even abuse raises serious safety concerns. Model developers have made tremendous efforts to introduce safety mechanisms that can address these concerns in T2I models. However, the existing safety mechanisms, whether external or internal, either remain susceptible to evasion under distribution shifts or require extensive model-specific adjustments. To address these limitations, we introduce Safe-Control, an innovative plug-and-play safety patch designed to mitigate unsafe content generation in T2I models. Using data-driven strategies and safety-aware conditions, Safe-Control injects safety control signals into the locked T2I model, acting as an “update” in a pacth-like manner. Model developers can also construct various safety patches to meet the evolving safety requirements, which can be flexibly merged into a single, unified patch. Its plug-and-play design further ensures adaptability, making it compatible with other T2I models of similar denoising architecture. We conduct extensive evaluations on six diverse and public T2I models. Empirical results highlight that Safe-Control is effective in reducing unsafe content generation across six diverse T2I models with similar generative architectures, yet it successfully maintains the quality and text alignment of benign images. Compared to seven state-of-the-art safety mechanisms, including both external and internal defenses, Safe-Control significantly outperforms all baselines in reducing unsafe content generation. For example, it reduces the probability of unsafe content generation to 7%, compared to approximately 20% for most baseline methods, under both unsafe prompts and the latest adversarial attacks.
Index Terms:
Text-to-Image, safety patch, defense mechanisms.
I Introduction
Text-to-image (T2I) generation models are emerging techniques that produce high-quality images based on natural language descriptions in the input prompt, such as “A photo of a hill full of sheep.” These models have quickly gained significant popularity due to their ease of use, along with the quality and flexibility of the generated images. Among them, the most representative variants of T2I model include Stable diffusion [1], Imagen [2], and DALL·E 3 [3]. These models are being widely adopted across various industries, significantly transforming sectors worldwide. For example, advertisers and marketers are leveraging this technology to create compelling content that enhances consumer engagement and brand recall.
Motivation. Despite the advancements in T2I models, their potential for misuse or even abuse raises serious safety concerns. Recent studies [4, 5] have demonstrated that the T2I models prone to generate NSFW (Not Safe For Work) imagery, such as those related to violence and child-unsafe, when prompted with unsafe text prompts. Quantifying and mitigating T2I models’ unsafe content generation thus become increasingly important research topics.
Figure 1: An illustration of our proposed Safe-Control. It locks down production-ready T2I generation models (the “system”) and applies conditional safety controls (the “patch”) to the model during its image generation process.
Current research in this field primarily focuses on preventing unsafe image generation through external or internal defenses. External defense methods [6, 7], such as safety filters, detect unsafe inputs or outputs during image generation. While efficient to deploy, these filters do not fundamentally block unsafe content and can be bypassed due to distribution shifts. Internal defenses include two main strategies: inference guidance [8, 4] and model fine-tuning [9, 10]. Inference guidance methods aim to steer image generation away from unsafe content during inference. However, they rely heavily on the model’s existing knowledge of unsafeness, limiting their effectiveness. Model fine-tuning methods adjust T2I model parameters to remove unsafe concepts. While effective at blocking explicit unsafe content, they can degrade model performance and are not transferable to other models, increasing implementation costs.
Contributions In this paper, we introduce Safe-Control, an innovative plug-and-play safety patch designed to mitigate unsafe content generation in T2I models. The concept of Safe-Control can be viewed as a system patch, with its update process resembling that of patch management in operating systems. System patches are typically used to fix vulnerabilities, improve performance, or add new features, while Safe-Control focuses on enhancing the safety of T2I models to reduce the unsafe content generation.
Specifically, Safe-Control functions by locking down production-ready T2I models (the “system”) and integrating an additional end-to-end model that applies conditional safety controls (the “patch”) during the image generation process. See Fig. 1 for an illustration. Using data-driven strategies and safety conditions, Safe-Control injects safety control signals into the T2I model in a patch-like manner, acting as an “update” to the model. Abstractly, Safe-Control incorporates four key design elements:
  • Model Parameters Preservation: Safe-Control locks the parameters of production-ready T2I models, ensuring that their ability to generate high-quality images remains minimally impacted.
  • Pretrained Encoding Utilization: Safe-Control reuses the trainable copy of the T2I model’s encoding layers. These layers (essentially a model), pretrained on billions of images, serve as a robust backbone, enabling Safe-Control to learn a diverse set of safety conditions’ visual semantics.
  • Conditional Safety Controls: Safe-Control creates a multi-modal dataset containing pairs of malicious prompts, corresponding safe images, and specific safety conditions from various unsafe domains. For instance, the prompt “a naked woman” would pair with an image of a clothed woman and the safety condition, “Add clothing to the person in the image.” By training on this dataset, Safe-Control learns to generate image-level safety control signals, making it more robust in suppressing unsafe content, regardless of the input prompt.
  • Zero Convolution Connection: Safe-Control is connected to the original, locked model via zero convolution layers, with weights initialized to zero. These weights gradually increase during training, allowing conditional safety control signals to be integrated into the deep features of the large T2I models.
Note that the trainable encoding model with learned safety control signals is considered a patch or “update” for the model. As safety requirements evolve, model developers can train additional safety patches to enhance their T2I models’s safety. These patches, both new and existing, can be merged using weighted techniques [11, 12, 13, 14] to create a unified safety patch. Besides, the plug-and-play design of Safe-Control also ensures significant transferability, allowing it to be easily adapted to other T2I models with similar generative architectures.
We have implemented a prototype of Safe-Control and conducted extensive evaluation on six diverse, public, and popular T2I models, such as Comic Diffusion [15], Deliberate [16], and PerfectWorld [17]. Empirical results highlight that Safe-Control is effective in reducing unsafe content generation across all T2I models, yet it successfully maintains the quality and text alignment of benign images. In comparison with seven state-of-the-art safety mechanisms, including both external defenses (e.g., Safety Checker [6]) and internal defenses (e.g., SLD [6] and SafeGEN [10]), Safe-Control significantly outperforms all baselines in mitigating unsafe content generation, both under natural malicious prompts and in response to the latest malicious attacks (e.g., SneakyPrompt [18] and Ring-A-Bell [19]). Specifically, the results under natural malicious prompts show Safe-Control reduces unsafe content probability to 7%, compared to around 20% for most baseline methods on I2P prompt datasets. Besides, Safe-Control maintains consistently low unsafe probability across two malicious attacks, demonstrating its robustness in protecting against adversarial threats. Importantly, Safe-Control outperforms the baselines in terms of CLIP scores, indicating that our proposed defense method better maintains the text alignment capability of the original T2I models. Lastly, we perform experiments to investigate how different hyperparameters impact Safe-Control. The results suggest that the size of the training data is inversely proportional to the number of iterations required by Safe-Control to achieve its optimal defensive performance.
In summary, our contributions are as follows:
  • New Defense. We propose a plug-and-play safety patch for T2I generation models, named Safe-Control, in a patch-like manner, which mitigates unsafe image generation while preserving the generative capabilities of original production-ready models.
  • Novel Technique. We propose a safety condition mapping module to handle non-visualizable conditional safety controls. Additionally, we adopt the design principles of model parameters preservation and zero convolution connection to preserve benign content generation of T2I generation models.
  • Extensive Evaluation. We conduct extensive evaluations of Safe-Control across six diverse T2I models with similar generative architectures, showcasing its effectiveness and transferability. Besides, our results demonstrate that Safe-Control significantly outperforms seven state-of-the-art baselines, both under natural malicious prompts and in response to the latest malicious attacks.
II Background
II-A Unsafe Content Generation in T2I Models
Text-to-image generation models have gained popularity due to their ease of use and high-quality, flexible images. However, Birhane et al. [20] raise concerns about datasets scraped from the internet, such as LAION-400M [21], which lack content moderation, potentially leading to unsafe content generation.
The definition of unsafe content varies by context and culture, making it subjective. In this paper, we focus on images containing hate, harassment, violence, self-harm, sexual content, shocking material, illegal activities, or nudity, as outlined in the OpenAI content policy [22] and Gebru et al. [23].
II-B Safety Mechanisms for T2I Models
Present strategies have two categories: external and internal defenses.
External Defenses. External defenses, like safety checker [6] officially released by SD, are efficient for deployment but suffer from under-generalization and vulnerability to adversarial prompts due to distribution shifts. Similarly, SD v2.1 [7] retrains on censored data with these filters, but this approach can be computationally expensive, may not fully remove harmful content, and could reduce model performance.
Internal Defenses. Internal defenses use two strategies: guiding generation to avoid unsafe content (e.g., SLD [4] and InterpretDiffusion [8]) or fine-tuning models to remove unsafe concepts (e.g., ESD [9] and SafeGEN [10]). The first relies on the model’s existing safety knowledge, limiting adaptability to new threats. The second risks “catastrophic forgetting” and lacks cross-model applicability. In contrast, our approach is model-agnostic and preserves production-ready models, ensuring robustness against adversarial prompts while maintaining performance for benign samples.
II-C Threat Model
We outline the goals and capabilities of the adversary and defender.
II-C1 Adversary
We first present the adversary’s goal and capability.
  • Goal: The adversary’s primary goal is to exploit a T2I model to generate content that violates safety or ethical standards, either deliberately or by evading any existing safeguards.
  • Capability: We assume that the adversary has closed-box access to an online text-to-image model and may query the model with prompts.
II-C2 Defender
Then we present the defender’s goal and capability in this paper.
  • Goal: The defender, primarily the model developer, has two key objectives: (1) to prevent the T2I model from generating unsafe content under any input prompt; and (2) to ensure that safety mechanisms have no adverse effect on benign content generation.
  • Capability: The defender is assumed to have full access to the model parameters and the ability to deploy safety mechanisms as needed.
III Design of Safe-Control
III-A Overview
Design Motivation. While safety mechanisms for T2I models exist, they lack robustness in reducing unsafe content, maintaining generative quality, and transferring across models. Inspired by ControlNet [24] in image editing, our key idea is to create an external safety patch—an end-to-end model—that injects safe control signals during generation without altering the target T2I model. Specifically, we lock the production-ready model and train the safety patch using unsafe prompts and safe image pairs, enabling it to remove or modify unsafe concepts, like adding clothes to generated images of naked people.
Figure 2: The overall architecture of Safe-Control, including the preliminary generation of training data and the training stage.
Challenges. The mechanism of Safe-Control is data-driven; however, the absence of publicly available datasets presents the first challenge: how to construct training data pairs of unsafe text and corresponding safe images. Besides, in image editing, ControlNet enables explicit spatial, structural, or geometric control over generated content using visual conditions (e.g., canny edges, human key points). However, applying this approach to T2I model safety is challenging, as no visual condition effectively represents “safety.” This leads to the second challenge: how to design safety conditions to ensure safe content generation in T2I models. Finally, preserving the generative capability of the original model is essential. Thus, the third challenge is how to train Safe-Control to avoid introducing harmful noise during the diffusion process for benign content generation.
Methodology Outline. To address these challenges, we propose Safe-Control, with an overview provided in Fig. 2.
The figure left side shows the Training Data Generation process, where we construct a multi-modal dataset by pairing unsafe prompts (e.g., “a naked female”) with corresponding safe images (see Section III-B). Safe prompts are generated from unsafe prompts using a Large Language Model (LLM) and a crafted prompt template, guided by the OpenAI content policy [22] and the definition of unsafe content from Gebru et al. [23]. These safe prompts are fed into a third-party text-to-image model to produce candidate images, which are then filtered using semantic tools and content classifiers to ensure safety. Training Safe-Control on this dataset allows it to learn conditional safety signals, enabling the transformation of unsafe images generated from unsafe prompts into safe ones.
The figure right side shows the Safe-Control training framework. Safe-Control adopts its network architecture from ControlNet [24], using a trainable copy of the target T2I model’s encoding layers as the backbone of the safety patch. Additionally, Safe-Control employs zero convolution connections between the trainable copy and the target T2I model. To enhance performance and safety, Safe-Control implements two key strategies: First, it locks the parameters of the target T2I model, training only the external safety patch. This preserves the model’s ability to generate high-quality images. Second, Safe-Control is trained on both a multi-modal dataset and a benign dataset containing safe prompts and images. This ensures that only conditional safety signals are injected, protecting the model from harmful noise without compromising its utility. Furthermore, Since visual conditions cannot effectively represent “safety,” directly applying ControlNet for safety enforcement is challenging. To address this, we replace ControlNet’s visual condition with safety modification instructions for input prompts, serving as the safety condition to guide the T2I model’s generation process. Additionally, we introduce a safety condition mapping module to handle non-visual safety conditions. See details in Section III-C.
TABLE I: Example pairs of unsafe prompts and their corresponding LLM-generated safe prompts.
Category     unsafe Prompt     Safe Prompt
Sexual     A naked woman posing in a hot springs with her legs crossed     A woman wearing a bathing suit posing in a hot spring with her legs crossed
    A nude woman with big breasts sitting on a black leather couch     A woman wearing a modest outfit is sitting on a black leather couch
Hate     Graphic of a woman shouting at a man aggressively     Graphic of a woman talking with a man
Please generate a drawing of a big group of sad children and toddlers with a symbol
of bleeding hearts on their chests, boldly depicted. display them in light and darkness
Please generate a drawing of a big group of upset children and toddlers with a symbol
of loving hearts on their chests, boldly depicted. display them in light and darkness.
III-B Training Data Generation
For the benign dataset, we select a subset from Flickr30k [25]. For the multi-modal dataset, which pairs unsafe prompts with safe images, we generate safe images instead of rejecting the prompts entirely. This ensures that users without harmful intent still receive appropriate content. For example, the prompt “A woman standing by the swimming pool” may seem harmless but could lead a T2I model to generate inappropriate content. We collect unsafe prompts from Lexica [^1] https://lexica.art/, ensuring no overlap with the evaluation dataset. A detailed description follows, and we will release the dataset later.
Figure 3: The prompt template for generating safe prompts corresponding to unsafe prompts in ChatGPT 3.5.
Generating Safe Prompts. To ensure reliable safe prompts, we use LLMs to generate multiple candidate prompts for each unsafe prompt rather than a single substitution. We define unsafe scenarios based on OpenAI’s content policy [22] and Gebru et al. [23], rather than relying solely on the LLM’s inherent understanding. This approach provides clear context guidance and standardizes unsafe content criteria across the dataset. Fig. 3 presents the prompt template, and Table I lists example prompt pairs.
Generating Safe Images. For each safe prompt, we use a T2I model (e.g., Stable Diffusion [26]) to generate multiple candidate images. We then apply semantic filtering tools, such as CLIPScore [27] and content classifiers (e.g., NudeNet [28]), to select the most accurate image while ensuring safety. To enhance precision and prevent misclassification, we raise classifier thresholds.
In addition, to mitigate potential biases introduced by LLMs, we implement a post-generation review process that leverages the OpenAI content policy [22] guidelines along with external content moderation tools, such as FairFace [29]. This process ensures that the resulting images not only adhere to safety standards but also reflect a balanced representation of context and intent.
III-C Finding Safety Condition
ControlNet is designed to regulate content generation based on explicit visual conditions such as posture or depth. However, “safety” is difficult to define as a unified visual condition. Instead, we leverage textual descriptions to enforce safety. For example, for an unsafe prompt like “a naked woman,” the modification instruction could be “Put clothes on the person in the image.” For a neutral prompt like “a monkey,” no modification is needed. Similarly, concepts related to hate or violence can be managed through specific instructions. We introduce safety modification instructions as a control mechanism for T2I models, replacing traditional visual conditions. To achieve this, we use an LLM to assess the prompt’s safety and generate corresponding modification instructions. The prompt templates follow those in Section III-B.
III-D Training Safe-Control
Suppose $\epsilon_{\theta_{g}}\left(\cdot\right)$ is the noise predictor network of the target T2I model, with parameters $\theta_{g}$, that estimates the current noise $\tilde{\epsilon}$ based on the given noisy image $z_{t}$ and the user-input prompt $c_{p}$ at time step $t$ as
$\tilde{\epsilon}=\epsilon_{\theta_{g}}\left(z_{t},t,c_{p}\right)$ (1)
where $c_{p}$ is the text encoding transformed by the text prompt $p$ using text encoder $\tau$ and $z_{t}$ is the noisy latent representation transformed by the pixel-space image $x$ using T2I model’s encoder $\omega$.
As aforementioned, Safe-Control creates a trainable copy of the model’s encoding layers, serving as a robust backbone. Pretrained on billions of images, these layers help Safe-Control quickly adapt to diverse safety conditions and generalize to unseen unsafe categories. Besides, fine-tuning these layers balances a strong semantic foundation with flexibility for task-specific safety requirements. Here, we denote the copied encoding layers and zero convolution connections as $\epsilon_{\theta_{e}^{\ast}}$, with parameters $\theta_{e}^{\ast}$.
Algorithm 1 Training Safe-Control
0: Data: $\mathcal{X}=\left(\left\{p_{i}^{m},x_{i}^{s}\right\}_{i=1}^{N}\cup\left\{p_{j}^{b},x_{j}^{b}\right\}_{j=1}^{M}\right)$, Noise level scheduler: $\left\{\beta_{1},...,\beta_{T}\right\}$, Safety condition: $s$, Text encoder: $\tau$, Image encoder: $\omega$
1:repeat
2:  $t\sim Uniform\left(\left\{1,...,T\right\}\right)$
3:  $\epsilon\sim\mathcal{N}\left(0,\mathrm{I}\right)$
4:  $z_{0}=\omega\left(x\right)$; $c^{o}=\tau\left(p\right)$; $c^{s}=\tau\left(s\right)$
5:  $z_{t}=\sqrt{\bar{\alpha_{t}}}z_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon$
6:  Take gradient descent step on
7:  $\qquad\nabla_{\theta_{e}^{\ast},\theta_{m}}\left\|\epsilon-\epsilon_{\theta_{g}\oplus\theta_{e}^{\ast}}\left(z_{t},t,c_{p},\mathcal{M}_{\theta_{m}}\left(z_{t},c_{s}\right)\right)\right\|_{2}^{2}$
8:until reach max epoch
[…]
Note that our approach introduces a safety condition mapping network, $\mathcal{M}_{\theta_{m}}$, to handle non-visualizable safety conditions. Specifically, to process text-based safety conditions, we use a cross-attention network to embed the safety condition $s$ into the noisy image. Given the encoded safety condition $c_{s}$ and the noisy image $z_{t}$ at time step $t$, the network outputs a safety “vision condition” for generating safety control signals, denoted as $\mathcal{M}_{\theta_{m}}(z_{t},c_{s})$.
To train Safe-Control, we first select data pairs <$p$, xx>from our dataset, including unsafe samples $p^{m}$, $x^{s}$>and benign samples $p^{b}$, $x^{b}$>. The T2I model’s encoder $\omega$ converts the image $x$ into a latent representation $z_{0}$, while the text encoder $\tau$ transforms the prompt $p$ into a text encoding $c_{p}$. We then use an LLM (e.g., ChatGPT) to determine the safety condition $s$ for each prompt and encode it as $c_{s}$. Finally, the DDPM noise scheduler [30] progressively adds noise $\epsilon$ to the image at a random time step $t$ based on a predefined schedule $\beta_{1},...,\beta_{T}$, generating a noisy image $z_{t}$ as follows:
$z_{t}=\sqrt{\bar{\alpha_{t}}}z_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon$ (2)
where $\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{i}=1-\beta_{i}$. The training loss is
$\mathcal{L}=\left\|\epsilon-\epsilon_{\left(\theta_{g},\theta_{e}^{\ast}\right)}\left(z_{t},t,c_{p},\mathcal{M}_{\theta_{m}}\left(z_{t},c_{s}\right)\right)\right\|_{2}^{2}$ (3)
where $\mathcal{L}$ is the overall learning objective of the entire generation model, which is directly inherited from ControlNet finetuning.
To avoid causing degradation of the original T2I model, we freeze the parameters $\theta_{g}$ of the original model and only optimize our proposed Safe-Control’s parameters $\theta_{g}^{\ast}$.
$\nabla_{\theta_{e}^{\ast},\theta_{m}}\left\|\epsilon-\epsilon_{\theta_{g}\oplus\theta_{e}^{\ast}}\left(z_{t},t,c_{p},\mathcal{M}_{\theta_{m}}\left(z_{t},c_{s}\right)\right)\right\|_{2}^{2}$ (4)
III-E Inference
Through the training process described in Algorithm 1, model developers can obtain a collection of safety patches (Safe-Control) for defending against different safety risks. For the new T2I model, model developers can directly add the outputs of the selected Safe-Control to the decoding layers of the T2I model for safe generation (see Fig. 2).
IV Experiment
Safe-Control is a flexible safety defense strategy that can be customized and trained by model developers or owners to meet the specific safety requirements of their T2I models. In this paper, we validate its effectiveness on the Stable Diffusion, which is one of the most popular generative model architectures. In particular, we consider various derivative models based on this generative model architecture. Our extensive experiments answer the following research questions (RQs):
  • RQ1. How effective is Safe-Control in mitigating the unsafe content generation across a variety of models with similar generative architectures?
  • RQ2. How does Safe-Control perform compared with different baselines?
  • RQ3. How robustness is Safe-Control against malicious attacks?
  • RQ4. How do different hyperparameters affect the performance of Safe-Control?
IV-A Experimental Settings
IV-A1 Datasets
We evaluate Safe-Control’s performance using the following four public prompt datasets. Specifically, we use three malicious prompt dataset, <country> body, I2P [4], and NSFW-200 [18] to evaluation Safe-Control’s effectiveness in mitigating unsafe content generation. Additionally, we employ a benign prompt dataset, MS COCO-2017 [31], to evaluate Safe-Control’s utility in preserving the high-fidelity of benign generation.
  • “<country> body”: The dataset consists of 50 prompts like “<country> body” with top-50 GDP countries. These prompts appear innocent but are likely to lead the T2I models to generate nude content.
  • NSFW-200: The dataset consists of 200 malicious prompts with NSFW content, which are generated using ChatGPT with GPT 3.5 under the guidance of a post on Reddit [32].
  • I2P: The dataset is a benchmark created for evaluating measures against unsafe content generation in T2I models, comprises 4,703 real-world, user-generated prompts that produce unsafe images, including hate, harassment, violence, self-harm, shocking, sexual, and illegal activity.
  • MS COCO 2017: MS COCO dataset is a large-scale labeled image dataset containing a vast number of images depicting humans and everyday objects. We use the validation set for evaluating image fidelity.
IV-A2 Baselines
We compare Safe-Control with seven latest safety countermeasures for T2I models. Specifically, we divide these baselines into the following groups:
  • N/A: In this group, the original Stable Diffusion (SD) serves as the control group without any defensive measures. Here we follow previous work [4, 9] to use the officially provided SD v1.4 [26] as the original SD.
  • External Defenses: In this group, they check the model’s inputs or outputs against known unsafe concepts though external classification models. Here, we use the officially released image-based safety checker [6] in SD to examine its performance in mitigating unsafe generation. Additionally, this group endeavors to censor the training data to remove problematic content entirely prior to the initiation of training. Here, we select the latest Stable Diffusion v2-1 [7], a model trained on a subset of LAION-5B [21] systematically filtered to exclude explicit pornographic material by employing the LAION-NSFW classifier [33], to be the representative of this group.
  • Internal Defenses: Internal defense methods can be broadly categorized into two classes: Inference guidance and Model fine-tuning. Inference guidance methods steer the denoising diffusion process away from unsafe areas during inference. In this case, we select the state-of-the-art method, SLD [4] (medium safety levels) and InterpretDiffusion [8], for the purpose of comparison. Model fine-tuning methods finetune a pre-trained model to erase unsafe concepts. In this case, we select the state-of-the-art method, ESD [9], and SafeGEN [10], for the purpose of comparison.
IV-A3 Text-To-Image Generation Models
Besides using SD v1.4 as a baseline, we also evaluate six derivative T2I models based on the Stable Diffusion architecture to assess the transferability of Safe-Control.
  • SD v1.5: The SD v1.5 [34] is a general Text-to-Image model parallel to SD v1.4. It was released by Runway ML, a partner of Stability AI, not by the Stability AI which is responsible for SD v1.4.
  • Comic Diffusion v2: The Comic Diffusion [15] is a fine-tuned model based on the Stable Diffusion architecture for generating Western comic-style art.
  • Protogen Diffusion x3.4: The Protogen [35] was warm-started with SD v1.5 and fine-tuned on various high-quality image datasets, and version x3.4 is closer to the effect of realistic visuals.
  • Deliberate v3: The Deliberate [16] is a model that leans towards realism and places a greater emphasis on detail.
  • PerfectWorld: The PerfectWorld [17] is a semi-realistic model, aiming to realize beautiful realistic faces using sensual endless bodies.
  • Realistic Vision v5.1: The Realistic Vision series of models are suitable for drawing realistic images of people and animals, and the latest version is Realistic Vision V5.1 [36].
IV-A4 Metrics
We utilize the following metrics to evaluate a T2I model’s ability in safe generation from two perspectives: Unsafe content Reduction and Benign content preservation.
  • Nudity Probability: We utilize NudeNet [^2] https://github.com/notAI-tech/NudeNet to focus on evaluating the effectiveness in moderating explicit content. In this paper, we only classify exposed genitalia, breasts, and buttocks as nude content. The threshold is set to 0.7 due to high false-positive rates.
  • Unsafe Probability: We follow previous work [4, 9] to combine two classifiers, Q16 classifier [^3] https://github.com/ml-research/Q16 and NudeNet, to automatically evaluate unsafe content generation. The Q16 classifier is able to detect a wide range of unsafe content in images. If not specified otherwise, an image is classified as unsafe if one or both of the classifiers output the respective label.
  • FID Score: The Frechet Inception Distance (FID) score [37] is a metric used to assess the quality of generated images. For our study, we use the COCO val 2017 dataset as the reference set, which provides prompts and corresponding real images. A lower score on the FID score means that the generated images have a higher fidelity.
  • LPIPS score: The Learned Perceptual Image Patch Similarity (LPIPS) score [38] is a widely used metric for evaluating the fidelity of generated images. We also use the COCO val 2017 dataset as the reference set. A lower LPIPS score means that the generated images have a higher fidelity.
  • CLIP Score: The CLIP Score [27] is a reference-free metric designed to evaluate the alignment between the prompt and the corresponding generated image. A higher score on the CLIP score means that the T2I model can faithfully reflect the user’s prompt by way of images.
IV-B RQ1: Effectiveness in Mitigating Unsafe Generation
In this research question, we assess Safe-Control’s mitigation effectiveness as a patch, applied to a range of models.
TABLE II: The nudity probability of images generated by various T2I models, both with and without our proposed $\textit{Safe-Control}_{nudity}$ patch. The values marked in blue indicate the ratio of nudity content erased by $\textit{Safe-Control}_{nudity}$ relative to the original T2I models. A decrement is represented by “-,” whereas “+” signifies an increment.
Method Nudity Probability $\downarrow$
<country> body I2P NSFW-200
SD-v1.4 0.22 0.17 0.36
SD-v1.4
+ $\textit{Safe-Control}_{nudity}$ 0.02 (-93.12%) 0.01 (-93.38%) 0.02 (-94.37%)
SD-v1.5 0.19 0.20 0.31
SD-v1.5
+ $\textit{Safe-Control}_{nudity}$ 0.01 (-93.78%) 0.01 (-93.06%) 0.00 (-100%)
Comic Diffusion v2 0.08 0.03 0.23
Comic Diffusion v2
+ $\textit{Safe-Control}_{nudity}$ 0.03 (-56.97%) 0.00 (-88.89%) 0.02 (-91.11%)
ProtogenDiffusion x3.4 0.38 0.14 0.46
ProtogenDiffusion x3.4
+ $\textit{Safe-Control}_{nudity}$ 0.03 (-91.98%) 0.05 (-66.40%) 0.13 (-72.83%)
Deliberate v3 0.67 0.38 0.74
Deliberate v3
+ $\textit{Safe-Control}_{nudity}$ 0.10 (-85.25%) 0.18 (-52.27%) 0.28 (-62.84%)
PerfectWorld 0.42 0.26 0.27
PerfectWorld
$\textit{Safe-Control}_{nudity}$ 0.08 (-79.91%) 0.07 (-71.93%) 0.09 (-67.92%)
RealisticVision v5.1 0.39 0.4 0.61
RealisticVision v5.1
+ $\textit{Safe-Control}_{nudity}$ 0.12 (-69.49%) 0.12 (-71.19%) 0.17 (-72.13%)
TABLE III: The unsafe probability of images generated by various T2I models, both with and without $\textit{Safe-Control}_{multiple}$ patch. The values marked in blue indicate the ratio of unsafe content erased by $\textit{Safe-Control}_{multiple}$ relative to the original T2I models. A decrement is represented by “-,” whereas “+” signifies an increment.
Method Unsafe Probability $\downarrow$
Sexual Self-harm Hate Violence Shocking Harassment Illegal activity Oveall
SD-v1.4 0.23 0.27 0.24 0.32 0.37 0.20 0.23 0.27
SD-v1.4 + $\textit{Safe-Control}_{multiple}$ 0.05 (-80%) 0.08 (-69%) 0.12 (-50%) 0.08 (-75%) 0.08 (-78%) 0.07 (-63%) 0.07 (-72%) 0.07 (-73%)
SD-v1.5 0.23 0.28 0.25 0.32 0.40 0.21 0.22 0.28
SD-v1.5 + $\textit{Safe-Control}_{multiple}$ 0.05 (-77%) 0.08 (-70%) 0.10 (-58%) 0.08 (-76%) 0.10 (-75%) 0.06 (-70%) 0.07 (-70%) 0.07 (-73%)
Comic Diffusion v2 0.06 0.09 0.17 0.12 0.18 0.11 0.10 0.11
Comic Diffusion v2 + $\textit{Safe-Control}_{multiple}$ 0.01 (-76%) 0.02 (-71%) 0.06 (-63%) 0.03 (-75%) 0.03 (-82%) 0.04 (-67%) 0.03 (-74%) 0.03 (-75%)
ProtogenDiffusion x3.4 0.18 0.28 0.27 0.25 0.39 0.16 0.20 0.25
ProtogenDiffusion x3.4 + $\textit{Safe-Control}_{multiple}$ 0.07 (-61%) 0.14 (-52%) 0.19 (-28%) 0.13 (-47%) 0.19 (-50%) 0.11 (-30%) 0.09 (-55%) 0.13 (-49%)
Deliberate v3 0.35 0.35 0.36 0.32 0.46 0.25 0.27 0.34
Deliberate v3 + $\textit{Safe-Control}_{multiple}$ 0.14 (-60%) 0.20 (-42%) 0.23 (-38%) 0.17 (-47%) 0.26 (-43%) 0.16 (-60%) 0.15 (-41%) 0.18 (-46%)
PerfectWorld 0.25 0.35 0.36 0.32 0.41 0.20 0.23 0.30
PerfectWorld + $\textit{Safe-Control}_{multiple}$ 0.07 (-71%) 0.15 (-57%) 0.16 (-55%) 0.14 (-57%) 0.19 (-52%) 0.10 (-47%) 0.09 (-60%) 0.13 (-57%)
RealisticVision v5.1 0.35 0.38 0.30 0.31 0.45 0.21 0.26 0.33
RealisticVision v5.1 $\textit{Safe-Control}_{multiple}$ 0.12 (-66%) 0.18 (-53%) 0.16 (-47%) 0.13 (-60%) 0.21 (-52%) 0.10 (-51%) 0.10 (-62%) 0.14 (-57%)
Average Erasing ratio     70% 59% 47% 61% 61% 51% 62%     61%
Setup. We train our proposed safety patch, Safe-Control, on SD v1.4 [26] and then directly transfer it as a safety patch onto six diverse T2I models with similar generative architectures. Subsequently, we utilize three prompt datasets, namely “<country> body”, I2P, and NSFW-200, to assess the effectiveness of Safe-Control in suppressing unsafe content generation. Additionally, we select a benign prompt dataset, namely MS COCO 2017, to evaluate Safe-Control’s ability of preserving the benign generation.
Nudity Content Reduction. Nudity, as one of the most severe categories within the ambit of unsafe content, poses substantial hazards due to its rampant spread. Consequently, this part will principally discuss the effectiveness of our proposed defensive methods Safe-Control in reducing nudity content.
To ensure the comprehensiveness of the evaluation, we selected three different types of prompt datasets. Table II shows the nudity probability of images generated by various T2I models, both with and without our proposed $\textit{Safe-Control}_{nudity}$ patch. Note that $\textit{Safe-Control}_{nudity}$ is a specifically trained safety patch designed to mitigate the generation of nude content. The blue-marked values in the table indicate the ratio of nudity content erased by $\textit{Safe-Control}_{nudity}$ relative to the original T2I models.
On the whole, $\textit{Safe-Control}_{nudity}$ has shown significant effects in suppressing the generation of nude content in T2I models. In the best case, it can eliminate over 93% of the explicit content in the original T2I models, and in the worst case, it can also eliminate more than half of the nude content. Specifically, given that $\textit{Safe-Control}_{nudity}$ was trained on the SD v1.4, this makes its ability to mitigate nudity from this model particularly noticeable. Across three distinct prompt datasets, our method is able to reduce over 93% of the nudity content present in the original model, effectively lowering the probability of nudity occurrence to nearly zero.
Furthermore, $\textit{Safe-Control}_{nudity}$ has demonstrated exceptional generalizability across various different models. For example, when $\textit{Safe-Control}_{nudity}$, which is trained on SD v1.4, is directly transferred to the other six T2I models, it can still, on average, reduce the generation of 77% of explicit content from the original models. In particular, this reduction can even exceed 93% on the SD v1.5, similar to the performance seen on the SD v1.4. The main reason is that SD v1.4 and SD v1.5 are initialized with the same weights of the SD v1.2 [39] checkpoint and subsequently fine-tuned on the same dataset, laion-aesthetics v2 5+  [^4] https://laion.ai/blog/laion-aesthetics/. In general, these findings indicate that Safe-Control is effective in reducing the generation of nudity-related content across various models with similar generative architectures, thereby potentially preventing the misuse of T2I models for nudity content creation.
Multiple Unsafe Contents Reduction. Apart from nudity, other forms of unsafe content, including hate, harassment, violence, self-harm, shocking, sexual, and illegal activity, also bear significant adverse societal impacts. In this part, we utilize the I2P prompt dataset (which includes all types of these unsafe contents) to validate the effectiveness and versatility of our proposed Safe-Control. In order to control the generation of multiple unsafe content simultaneously, we combined safety patches capable of suppressing different kinds of unsafe risks (such as, $\textit{Safe-Control}_{shocking}$ and $\textit{Safe-Control}_{sexual}$) using weighted techniques [11, 12, 13, 14] to create a unified safety patch, named $\textit{Safe-Control}_{multiple}$.
Table III delineates the unsafe probability of images generated by various T2I models, both with and without the unified safety patch $\textit{Safe-Control}_{multiple}$. The values marked in blue indicate the ratio of unsafe content erased by $\textit{Safe-Control}_{multiple}$ relative to the original T2I models. We have three observations. First, similarly to the observations made on the nudity content reduction, one can observe in Table III that the probability of all types of unsafe contents markedly decreases when enhanced with our proposed safety patch. For example, for any type of unsafe content, the average unsafe probability reduction achieved by $\textit{Safe-Control}_{multiple}$ across seven T2I models is close to or exceeds 50%, with the highest even reaching up to 70%. Second, $\textit{Safe-Control}_{multiple}$ demonstrates potent defensive capability across seven types of T2I models, managing to remove an average of 61% of unsafe content, thus showing prominent transferability. Third, $\textit{Safe-Control}_{multiple}$ excels at mitigating the generation of sexual content, being able to reduce the generation probability of such unsafe content to no more than 7% in five models. This is primarily because the harm caused by sexual content is relatively high. Therefore, in setting up safety patch combinations, we adopted a strategy of granting $\textit{Safe-Control}_{multiple}$ a higher weight to suppress the generation of such unsafe content. For practical applications, users can adjust the weight of each patch according to their needs, thereby designing their own safety patch.
TABLE IV: The performance of $\textit{Safe-Control}_{nudity}$ in maintaining benign generation for prompts from COCO 2017 val, and its comparison with the original T2I models.
Method COCO 2017 Val
CLIP Score $\uparrow$ LPIPS Score $\downarrow$ FID Score $\downarrow$
SD-v1.4 31.28 0.7562 25.22
SD-v1.4
+ $\textit{Safe-Control}_{nudity}$ 31.46 0.7562 24.90
SD-v1.5 31.32 0.7555 25.04
SD-v1.5
+ $\textit{Safe-Control}_{nudity}$ 31.16 0.7565 25.71
Comic Diffusion v2 30.70 0.7498 26.88
Comic Diffusion v2
+ $\textit{Safe-Control}_{nudity}$ 30.27 0.7589 29.04
ProtogenDiffusion x3.4 31.83 0.7486 28.47
ProtogenDiffusion x3.4
+ $\textit{Safe-Control}_{nudity}$ 31.69 0.7561 30.91
Deliberate v3 31.79 0.7404 30.17
Deliberate v3
+ $\textit{Safe-Control}_{nudity}$ 31.73 0.7364 28.19
PerfectWorld 31.33 0.7471 28.64
PerfectWorld
+ $\textit{Safe-Control}_{nudity}$ 31.54 0.7537 31.54
RealisticVision v5.1 31.80 0.7408 29.11
RealisticVision v5.1
+ $\textit{Safe-Control}_{nudity}$ 31.54 0.7413 25.89
Benign Generation Preservation. To evaluate the ability of preserving benign generation, we discuss the overall impact of Safe-Control on image fidelity and text alignment of original T2I models. Ideally, the approach should have no adverse effect on either, especially on benign images. In line with previous research [10], we employ a benign prompt dataset, MS COCO 2017, and report the FID, LPIPS, and CLIP scores of images generated by various T2I models, both with and without $\textit{Safe-Control}_{nudity}$ patch in Table IV. We obtain three key observations from Table IV.
First, $\textit{Safe-Control}_{nudity}$ achieves a CLIP score on par with the original T2I models, indicating that it can fully maintain the text alignment ability of the original T2I models. This result can be well explained by Fig. 7. The first two columns show that for malicious prompt input, Safe-Control transforms unsafe semantics into safe semantics while keeping other safe semantics unchanged as much as possible. The last two columns show that for benign prompt input, Safe-Control will not change the safe semantics, thereby maintaining the text alignment capability of the original T2I models.
Second, on SD v1.4, $\textit{Safe-Control}_{nudity}$ has no adverse effect on benign generation. In contrast, some metrics become more better, especially the FID score. This is mainly attributed to mixing a certain proportion of benign samples from another dataset into the training dataset and reducing the modification of benign samples through Safety condition control.
Third, when transferred to other T2I models, $\textit{Safe-Control}_{nudity}$ performs excellently in text alignment. Although there are slight variations in image quality across different models, according to [4], these changes do not affect human visual perception.
Figure 4: The results of $\textit{Safe-Control}_{nudity}$ and baselines in reducing the generation of various exposed body parts, where ‘M’ stands for male and ‘F’ stands for female.
TABLE V: The nudity removal performance of our proposed $\textit{Safe-Control}_{nudity}$ and baselines compared to original Text-to-Image generative models across different adversarial prompt datasets.
Method Numbers of nudity detected $\downarrow$ Nudity Removal Rate $\uparrow$
<country> body I2P NSFW_200 Total
Original SD 218 151 71 N/A
Safety Filter 59 38 33 67%
SD 2.1 70 89 20 68%
SLD 81 70 50 49%
InterpreteDiffusion 62 78 38 56%
ESD 18 15 2 93%
SAFEGEN 19 40 1 88%
$\textit{Safe-Control}_{nudity}$     15 10 4     94%
IV-C RQ2: Performance Comparison with Baselines
We then compare Safe-Control and baselines in mitigating unsafe content generation across different malicious prompt datasets.
Setup. We follow previous work [4, 9] to use the officially provided SD v1.4 [26] as the original SD. That is, Safe-Control and baselines will be deployed on SD v1.4 to suppress the generation of unsafe content. For the fine-tuning baselines, we adopt the pre-trained models provided by the official source. For additional types of baselines, we execute operations as per the configurations dictated by the original source. As for Safe-Control, we train it on SD v1.4 using 10k data pairs.
Nudity Content Reduction. Nudity is among the most serious categories of unsafe content. In this part, we focus on comparing the performance of Safe-Control and baselines in reducing nudity content generation.
Table V shows the performance of our proposed defense and baselines on nudity content reduction across three different prompt datasets. With respect to the “<country > body” prompt dataset, we produced an aggregate of 2000 images using the prompt “<country > body” associated with the top-50 GDP countries (20 images for each country). Based on the results across three different prompt datasets, $\textit{Safe-Control}_{nudity}$ has a more significant effect in mitigating nudity content generation than baselines, with the highest average nudity removal rate of 94%. In baselines, only the ESD method is comparable to ours in effectiveness, while the average nudity removal rate of other baselines is over 6% lower than ours.
In addition, we found that the effectiveness of the SLD method was not ideal, primarily because we chose the SLD-medium configuration. When using the stronger SLD-strong and SLD-max configurations, the average nudity removal rates reached 73% and 85%, respectively. However, these two configurations significantly degraded the quality and text alignment of benign generation, as shown in Table VII.
Additionally, Fig. 4 shows the performance of our proposed method and baselines in reducing the generation of various exposed body parts, e.g., M-Breasts or F-Breasts, on the I2P dataset, where ‘M’ stands for male and ‘F’ stands for female. We observed that $\textit{Safe-Control}_{nudity}$ achieved an average nudity removal rate of 97.6% for sensitive body parts (including female breasts, male breasts, female genitals, male genitals, and buttocks), outperforming all baseline methods. In contrast, for body parts with safe semantics, such as feet, $\textit{Safe-Control}_{nudity}$ does not make significant alterations, in order to minimize the impact on the original T2I generative model’s capabilities.
Multiple Unsafe Content Reduction. In this part, we extend the defense targets from a single to multiple unsafe content generation to evaluate the comprehensiveness of Safe-Control and baseline methods. We employ the I2P prompt dataset, which contains seven types of unsafe semantics such as hate, harassment, violence, self-harm, shocking, sexual, and illegal activities, as the target for our defense. The goal of Safe-Control and baselines is to effectively reduce the generation of above unsafe content when applied to original T2I models.
Table VI delineates the performance of our proposed $\textit{Safe-Control}_{multiple}$ and baselines on multiple unsafe content reduction on I2P prompt dataset. First, overall, $\textit{Safe-Control}_{multiple}$ almost reduced the probability of generating unsafe content from all types of unsafe prompts to below 10%, representing an overall improvement of 4% compared to the baselines. Second, due to the potential impact of distributional shifts on the filters used during in data censorship or inference-stage filtering, external defenses (such as SD v2.1 and safety filter) can only reduce the probability of generating unsafe content in T2I models by less than 20%, indicating that their generalization capabilities are insufficient.
In addition, we observed that SLD also demonstrates significant performance in suppressing various types of unsafe content, with even the medium safety level configuration reducing the overall probability of unsafe content to 12%. Therefore, we further investigated the performance comparison between higher safety level configurations of SLD and Safe-Control in suppressing the generation of multiple unsafe contents, as shown in Fig. 5.
The figure shows the relationship between image quality (FID/CLIP/LPIPS) on benign samples and the probability of generating unsafe content for $\textit{Safe-Control}_{multiple}$ and SLD under different safety level configurations. We can observe that as the image quality degrades (i.e., FID/LPIPS becomes larger and CLIP becomes smaller), two methods can significantly reduce the probability of generating unsafe content. However, among baselines with similar image quality (similar FID/CLIP/LPIPS), our $\textit{Safe-Control}_{multiple}$ consistently achieve better effectiveness (lower unsafe probability). Given that defensive methods must be practical without affecting the generation of benign content, our proposed Safe-Control outperforms the SLD methods.
TABLE VI: The performance of our proposed $\textit{Safe-Control}_{multiple}$ and baselines on multiple unsafe content reduction on I2P prompt dataset.
Method
Unsafe Probability ($\downarrow$)
    Sexual Self-harm Hate Violence Shocking Harassment Illegal activity     Oveall
Original SD     23% 27% 23% 32% 37% 20% 23% 27%
Safety Filter     8% 24% 18% 28% 31% 15% 22% 21%
SD 2.1     15% 27% 25% 27% 35% 22% 20% 24%
SLD     10% 10% 10% 14% 20% 11% 8% 12%
InterpreteDiffusion     10% 18% 23% 21% 29% 16% 15% 18%
ESD     10% 20% 18% 26% 27% 18% 19% 20%
SAFEGEN     7% 10% 13% 13% 18% 9% 7% 11%
$\textit{Safe-Control}_{multiple}$     5% 8% 12% 8% 8% 7% 7%     7%
Figure 5: Comparision between our proposed Safe-Control and SLD on I2P prompt dataset. The different points of each method represent different safety level configurations. The “h1” to “h5” are the different weight combination strategies of multiple Safe-Control, with safety levels ranging from low to high.
Benign Generation Preservation. In line with previous research [10], we employ a benign prompt dataset, MS COCO 2017, and report the FID, LPIPS, and CLIP scores of images generated by T2I models deployed various safety mechanisms.
Table VII displays the performance of $\textit{Safe-Control}_{multiple}$ in maintaining benign generation for prompts from MS COCO 2017 val, and its comparison with baselines. Note that the downward arrow $\downarrow$ indicates that a lower score is more desirable, while the upward arrow $\uparrow$ suggests that a higher score is preferable. We obtain two key observations from Table IV.
First, our proposed single safety patch (such as $\textit{Safe-Control}_{nudity}$) outperforms the baselines in terms of CLIP scores, indicating that our proposed defense method better maintains the text alignment capability of the original T2I models. Additionally, the FID and LPIPS metrics also obtain comparable performance with other baselines and the original Text-to-Image generation model. Second, compared to a single safety patch (e.g., $\textit{Safe-Control}_{nudity}$), the multi-category combination of Safe-Control introduces a richer set of safety control signals, thereby substantially enhancing the overall defense effectiveness. Nonetheless, this improvement comes at the cost of a moderate decline in image quality.
Nevertheless, the degradation in image quality remains within an acceptable range when compared to the baseline models. And users can adjust the combination weights of multiple security patches according to their needs to balance the trade-off between defensive performance and image quality degradation. Moreover, our proposed Safe-Control can be plugged in or out at any time without affecting the production-ready T2I models.
TABLE VII: The performance of $\textit{Safe-Control}_{multiple}$ in maintaining benign generation for prompts from COCO 2017 val, and its comparison with baselines. Note that the downward arrow $\downarrow$ indicates that a lower score is more desirable, while the upward arrow $\uparrow$ suggests that a higher score is preferable.
Method     COCO 2017 Val
    CLIP Score $\uparrow$     LPIPS Score $\downarrow$     FID $\downarrow$
Original SD     31.28     0.7562     25.22
Safety Filter     30.21     0.7569     25.99
SD 2.1     31.47     0.7465     24.01
SLD-Max     29.10     0.7699     36.46
SLD-Strong     29.91     0.7636     33.01
SLD-Medium     29.92     0.7634     32.75
SLD-Weak     31.23     0.7564     26.68
InterpreteDiffusion     31.00     0.7612     26.73
ESD     30.41     0.7574     24.58
SAFEGEN     30.61     0.7641     29.96
$\textit{Safe-Control}_{nudity}$     31.46     0.7562     24.90
$\textit{Safe-Control}_{multiple}$     30.54     0.7521     26.15
IV-D RQ3: Robustness Against Malicious Attacks
In this section, we adopt two latest malicious attacks to evaluate the robustness of our proposed Safe-Control and baselines.
Setup. Recent studies [18, 19, 40, 41] have proposed malicious attacks to test and challenge the safety of T2I models. To evaluate our defensive method, we examine two state-of-the-art attacks: SneakyPrompt [18] and Ring-A-Bell [19]. Our defense strategy aims to strengthen the T2I model’s inherent safety; thus, we modify SneakyPrompt’s objective to focus on inducing unsafe content generation rather than bypassing the Safety Filter. Additionally, we use Q16 [42] to assess image safety, updating SneakyPrompt’s Safety Filter to Q16 while keeping all other configurations as specified in the original paper.
Results. Table VIII shows the robustness evaluation of our proposed $\textit{Safe-Control}_{multiple}$ method compared to several baseline approaches—Original SD, SLD, InterpreteDiffusion, ESD, and SAFEGEN—under two malicious attacks. The evaluation metrics assess the probability of unsafe content in images generated by T2I models equipped with different defenses.
Under the SneakyPrompt attack, $\textit{Safe-Control}_{multiple}$ achieves a notably lower unsafe probability (9%) compared to Original SD (55%), highlighting its robustness. Other baselines, such as ESD (36%), SAFEGEN (57%), and InterpreteDiffusion (30%), show some resilience but are less effective than $\textit{Safe-Control}_{multiple}$. Prior work [4] notes that the Q16 classifier is relatively conservative, often categorizing safe images as inappropriate, which may explain why $\textit{Safe-Control}_{multiple}$ has a slightly higher unsafe probability than SLD (3%). A manual analysis further indicates that both SLD and $\textit{Safe-Control}_{multiple}$ reduce the unsafe probability of generated images under the SneakyPrompt attack to 0%.
Under the Ring-A-Bell attack, which tests for violent and nude content, $\textit{Safe-Control}_{multiple}$ demonstrates strong safety control, with unsafe probabilities of 24% for violence and 6% for nudity. This is a significant improvement over Original SD, which records 96% and 81%, respectively, and it also outperforms other baselines, including ESD (86% for violence, 21% for nudity) and SAFEGEN (91% for violence, 64% for nudity). These results highlight the robustness of $\textit{Safe-Control}_{multiple}$ in reducing unsafe content generation under adversarial conditions. These findings confirm that Safe-Control is an effective approach for enhancing the safety of T2I generation models, providing robust protection against unsafe content generation in adversarial contexts.
TABLE VIII: The robustness of our proposed $\textit{Safe-Control}_{multiple}$ and baselines against latest malicious attacks. We report the unsafe probability of images being generated by two attacks.
Method     SneakyPrompt [18]     Ring-A-Bell [19]
        violence nudity
Original SD     55%     96% 81%
SLD     3%     70% 82%
InterpreteDiffusion     30%     90% 80%
ESD     36%     86% 21%
SAFEGEN     57%     91% 24%
$\textit{Safe-Control}_{multiple}$     9%     24% 6%
IV-E RQ4: Exploration on Hyperparameters
In this section, we will discuss the impact of different hyperparameters on Safe-Control.
Different Sizes of Training Dataset. Fig. 6 shows the impact of different training dataset sizes (1K/5K/10K) on Safe-Control. These three sizes of training datasets consist of malicious prompt related to sexual and their corresponding safe images generating by our proposed method. The results indicate that as the size of training data increases, the time it takes for Safe-Control to achieve optimal defensive performance is reduced. For example, when the size of the training dataset is 1K, safe-control requires a duration of 50,000 iterations to reach optimal defensive performance. However, when the training dataset expands to 10K, it only needs 25,000 iteration duration. Additionally, with the size of training data increases, the defensive performance that Safe-Control can achieve gradually improves. Lastly, As the volume of training data increases, the defensive performance that safe-control can achieve gradually improves.
Different Iterations of Training. Fig. 6 presents the correlation between the iterations and defensive performance during the training process of Safe-Control. Based on the results from three sizes of training data, we found that before performing 10,000 iterations, the defensive performance of Safe-Control significantly improves. However, after completing 10,000 iterations, the enhancement of Safe-Control’s defensive performance slows down, and most of the time its changes fluctuate within a small range. In addition, under the premise of the same batch size, to achieve the same defensive performance, the larger the training dataset, the fewer iterations needed.
Figure 6: The relationship between the dataset size (1K/5K/10K) and iterations during the training process.
V Discussion and Conclusion
In this paper, we introduce Safe-Control, an innovative plug-and-play safety patch designed to mitigate unsafe content generation in T2I models. Safe-Control functions by locking down production-ready T2I models (the “system”) and integrating an additional end-to-end model that applies conditional safety controls (the “patch”) during the image generation process. The empirical results highlight that Safe-Control is effective in reducing unsafe content generation across six diverse T2I models with similar generative architectures, yet it successfully maintains the quality and text alignment of benign images. Furthermore, in comparison with seven state-of-the-art safety mechanisms, Safe-Control significantly outperforms all baselines in mitigating unsafe content generation, both under unsafe prompts and in response to the latest adversarial attacks.
Limitation Due to the limitations in foundational models’ ability to understand text and images, there is still significant room for improvement in the quality of the training datasets we have constructed. However, the proposed approach used to construct training datasets are still effective. Therefore, as the capabilities of foundational models improve, the quality of our constructed training datasets will also correspondingly increase, further enhancing the security defense performance of Safe-Control. Lastly, the assessment of unsafe content relies on NudeNet and Q16, which may introduce some measurement inconsistencies across experiments. However, our manual validation helped address this issue, confirming that the results align with the relative effectiveness of different mitigation strategies.
Impact Statements
In this work, we have constructed a dataset containing extensive unsafe prompts (e.g., sexual, hate, and violence) sourced from public online platforms. Due to ethical considerations, these samples are available upon request to researchers for research purposes only, contingent on institutional approvals. Additionally, all experiments are conducted using publicly available datasets and standard model architectures widely adopted in T2I research. To ensure ethical compliance, we have implemented measures such as masking when presenting generated images that may contain unsafe content.